OpenStudy (anonymous):

suppose that integral a to b f(x)dx = 7 find integral b to a sqrt(3)f(x)dx.

OpenStudy (anonymous):

\[ \int\limits_{a}^{b}f(x)dx =7 (what is)\int\limits_{b}^{a} \sqrt{3}f(x)dx\]

OpenStudy (anonymous):

ahhhhhhhhhhhhhhh...

OpenStudy (anonymous):

i am thinking it's easier than it looks and is maybe -7sqrt(3) but i really am not sure... i never know if i am doing any of this right!

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Promise? :)

OpenStudy (anonymous):

pinky promise XD

OpenStudy (anonymous):

Or we can spit on our hands and shake hand XDD